OpenStudy (anonymous):

how do u answer S = 4v2 for v (the two is an exponent)

OpenStudy (anonymous):

? i just need to know how to write it because i know the answer but i dont know exactly how to write it

OpenStudy (anonymous):

yea do u know how?

OpenStudy (amistre64):

Fefi, do you know what undoes a ^2 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh the radical sign

OpenStudy (amistre64):

hmmm ..... lets try this: 3^2 = 9 right? how do you get from 9 back to 3?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

S = 4v^2 S/4 = v^2 and you got the rest right?

OpenStudy (anonymous):

ok so the answer would be the radical of s/4 or s/2?

OpenStudy (amistre64):

sqrt(S/4) = sqrt(S)/2

OpenStudy (amistre64):

=v :)

OpenStudy (anonymous):

awesome thank you soooo much

OpenStudy (amistre64):

yw, and good luck

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

sqrt(S/4) = sqrt(S)/2